OpenStudy (anonymous):

Use the elimination method to solve the system of equations.( My number three key went out lol) 8u+three v =-17 three u +v=-5 The solution is ?

OpenStudy (anonymous):

8u+3v=-17 3u+v=-5 What could you multiply one of the equations by that would make it where when you add them together, one of the variables would be eliminated?

OpenStudy (anonymous):

multiplly the second equation by three

OpenStudy (anonymous):

negative three

OpenStudy (anonymous):

Yep. What do you get when you do that?

OpenStudy (anonymous):

-6u-three v=-15

OpenStudy (anonymous):

You shouldve got -9u-3v=15

OpenStudy (anonymous):

8u+ three v=-17 -6u= three v=15

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

yeah i see

OpenStudy (anonymous):

Yep now add the equations 8u+3v=-17 -9u-3v=15 what do you get?

OpenStudy (anonymous):

u=2

OpenStudy (anonymous):

i subtract -9u from 8

OpenStudy (anonymous):

and -17 from 15

OpenStudy (anonymous):

Yep, so now that you know u=2 plug that in for u in one of the original equations and solve for v.

OpenStudy (anonymous):

well it says substitute

OpenStudy (anonymous):

so 2(2)+v=-5

OpenStudy (anonymous):

3(2)+v=-5

OpenStudy (anonymous):

ok so v=-1

OpenStudy (anonymous):

Nope, v=-11 v=-5-6 v=-11

OpenStudy (anonymous):

6+5

OpenStudy (anonymous):

not 6-5

OpenStudy (anonymous):

3(2)+v=-5 6+v=-5 subtract 6 from both sides v=-11

OpenStudy (anonymous):

so the solution is (2,-11)

OpenStudy (anonymous):

Yep. u=2 v=-11

OpenStudy (anonymous):

ok thanks could you help me with another problem?

OpenStudy (anonymous):

I could try

OpenStudy (anonymous):

lol well thats all we could do

OpenStudy (anonymous):

i am typing problem

OpenStudy (anonymous):

solve by elimination method.(my number three key is out still tryna fix...) 5a+2b=-three9 -5a+b=48 the solution is?

OpenStudy (anonymous):

lol, this one already has one that cancels out, so just go ahead and add them.

OpenStudy (anonymous):

please show me what you are saying

OpenStudy (anonymous):

ook

OpenStudy (anonymous):

b+5a=-three9

OpenStudy (anonymous):

b+5a=48 5a+2b=-three9

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

ok you added the two end number 48 and three9

OpenStudy (anonymous):

2b+b=three b

OpenStudy (anonymous):

Did you type the equations wrong the first time?

OpenStudy (anonymous):

no just whn i attempt i did

OpenStudy (anonymous):

First you said 5a+2b=-three9 -5a+b=48 Then you said b+5a=48 5a+2b=-three9 so which one are we doing?

OpenStudy (anonymous):

th first one is correct

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

so i have to divide

OpenStudy (anonymous):

29

OpenStudy (anonymous):

b=29

OpenStudy (anonymous):

what about solving for a

OpenStudy (anonymous):

-5a+(29)=48

OpenStudy (anonymous):

i think thats correct

OpenStudy (anonymous):

ok i do not know what the hck im doing on this but how i seet thee problem up for wwhat it says that what it is but all that i tried to attempt to do i do know if im right or wrong

OpenStudy (anonymous):

i dont know if im right or wrong

OpenStudy (anonymous):

What are your answers?

OpenStudy (anonymous):

im sorry if i confusd u

OpenStudy (anonymous):

(29, ?)

OpenStudy (anonymous):

i dont know how to do thee second step

OpenStudy (anonymous):

Lets start fresh. The equations are: 5a+2b=-39 -5a+b=48 Add them and you get 3b=9 b=3 Now that you know b, substitute it in and solve for a. 5a+2(3)=-39 5a=-45 a=-9

OpenStudy (anonymous):

ok i will

OpenStudy (anonymous):

So a=-9 b=3

OpenStudy (anonymous):

so it is -9, 1/three

OpenStudy (anonymous):

solution